OpenStudy (anonymous):

Solve x + 7 = –2(4x + 10). show all work please

OpenStudy (lgbasallote):

distribute -2 into 4x + 10 first x + 7 = -8x - 20 can you solve it now?

OpenStudy (anonymous):

x + 8x = -20 - 7 9x = -27 x = -27/9 x = -3